OpenStudy (anonymous):

Laplace solving system dx/dt = -x +y dy/dt = 2x x(0)=0 y(0)=1

OpenStudy (anonymous):

pew, last question

OpenStudy (accessdenied):

Hm. Did you know how to go about starting this one?

OpenStudy (anonymous):

I do, but am not sure what to do after getting x(s) and Y(s)

OpenStudy (anonymous):

X(s)(s+1)=Y(s) 2x+1=sY(s)

OpenStudy (accessdenied):

I believe you are just looking at solving for each like it were a linear system of equations. The first equation could be substituted into the second to let you find X(s), then find Y(s) using the result of X(s)...

OpenStudy (anonymous):

X(s)(s+1)=Y(s) 2X(s)+1=sY(s)

OpenStudy (accessdenied):

Y(s) = X(s) (s + 1) --> 2 X(s) + 1 = s Y(s) => 2 X(s) + 1 = s (X(s) (s + 1)) <-- this is a one variable system ignoring the s, so we can isolate the X(s).

OpenStudy (anonymous):

that's it? heh, that's not too bad. hmmn I can't seem to remember how to easily get Y(s) after though

OpenStudy (anonymous):

just plug it in? I suppose

OpenStudy (accessdenied):

Yeah, whatever you get from X(s), plug that back into either of the two equations to solve for Y(s). The first is easier in this case.

OpenStudy (anonymous):

pew, that was easy. Thanks, outta get some sleep lol

OpenStudy (accessdenied):

Haha, yea. I should be going to bed also... :P Good job, and gnight!

OpenStudy (anonymous):

thanks again for the help!